How do you do this


7. A ball is thrown vertically downward from the edge of a bridge 25.5 m high with an initial speed of 18.0 m/s. The ball falls all the way down and strikes the water below. Determine the time it takes the ball to strike the water
Answer #1

From

Y=Y_{o}+V_{oy}t+\frac{1}{2}a_{y}t^{2}

When the ball hits water Y=0

=>0=25.5+18t+(1/2)(-9.81)t2

4.905t2-18t-25.5=0

solving quadratic equation we get

t=4.76 s